# Additive Complementary Pairs of Codes

### v2

@@ Abstract
An additive code is an ${\rm I\!F}_{q}$-linear subspace of ${\rm I\!F}_{q^{m}}^{n}$ over ${\rm I\!F}_{q^{m}}$, which is not a linear subspace over ${\rm I\!F}_{q^{m}}$. Linear complementary pairs(LCP) of codes have important roles in cryptography, such as increasing the speed and capacity of digital communication and strengthening security by improving the encryption necessities to resist cryptanalytic attacks. This paper studies an algebraic structure of additive complementary pairs (ACP) of codes over ${\rm I\!F}_{q^{m}}$. Further, we characterize an ACP of codes in analogous generator matrices and parity check matrices. Additionally, we identify a necessary condition for an ACP of codes. Besides, we present some constructions of an ACP of codes over ${\rm I\!F}_{q^{m}}$ from LCP codes over ${\rm I\!F}_{q^{m}}$ and also from an LCP of codes over ${\rm I\!F}_{q}$. Finally, we study the constacyclic ACP of codes over ${\rm I\!F}_{q^{m}}$ and the counting of the constacyclic ACP of codes. As an application of our study, we consider a class of quantum codes called Entanglement Assisted Quantum Error Correcting Code (EAQEC codes). As a consequence, we derive some EAQEC codes.
@@ 1 Introduction
Linear complementary pairs (LCP) of codes, which were introduced by Bhasin et al. in [22], are extensively explored for wide application in cryptography (see [2, 6]). Further, we refer to some papers [1, 7, 11] on the LCP of codes over a finite field, and we point out that complementary pairs of codes without linearity properties have an essential role in constructing quantum code (see [23]). Thus, it is a useful context for studying non-linear complementary pairs of codes in coding theory. In particular, we study additive complementary pairs (ACP) of codes that extend the analogous of additive complementary dual (ACD) codes. With this point of view, lots of research has been done for additive complementary dual (ACD) codes (see [24, 25, 26]). LCP of codes is a generalization of linear complementary dual (LCD) codes. The notion of LCD codes was introduced by Massy in [20] and later studied in [8, 9, 10, 11, 21].
@@
Besides, the study of additive codes has gathered significant attention due to their theoretical significance and practical applications. One prominent application of additive codes lies in quantum error correction, where they are employed to protect quantum information from errors induced by noise or other environmental factors. Additionally, additive codes find utility in secret-sharing schemes, where they facilitate the secure distribution of confidential information among multiple parties, ensuring that only authorized subsets of participants can reconstruct the original secret [18]. Additive codes represent a crucial class of codes within coding theory, introduced by Delsarte et al. [13]. Generally, additive codes are subgroups of the underlying abelian group. Further, Huffman [17] provided an algebraic structure for additive cyclic codes over ${\rm I\!F}_{4}$. Through ongoing research and exploration, the potential of additive codes continues to expand, offering valuable insights and solutions in the realms of information theory, quantum computing, and cryptography (see [14]). Recently, Shi et al. [25] developed a theory of ACD codes over ${\rm I\!F}_{4}$ for trace Euclidean and Hermitian inner products. The authors introduced a nice construction of ACP codes over ${\rm I\!F}_{4}$ from binary codes in the same paper. They also established that ACD codes are potentially related to LCD codes. In the same spirit, Choi et al. [12] studied ACP codes over finite field ${\rm I\!F}_{q^{m}}$ with $m\geq 2$ in a general setup of trace inner product. They also provided new construction of ACP codes with respect to trace-Euclidean, Hermitian, and Galois inner products. This general setup showed that ACD codes are related to LCD codes. In this same paper, they further computed good numerical examples of ACP codes. Motivated by these papers, we study the theory of additive complementary pair (ACP) of codes. We further derive some construction of ACP of codes. This paper establishes a relationship between the LCP of codes and the ACP of codes.
@@
On the other hand, an $\dfrac{{\rm I\!F}_{q}[X]}{\left(X^{n}-\lambda\right)}$-submodule of $\dfrac{{\rm I\!F}_{q^{m}}[X]}{\left(X^{n}-\lambda\right)}$ is a $\lambda$-constacyclic ${\rm I\!F}_{q}$-linear additive code over ${\rm I\!F}_{q^{m}}$, where $\lambda\in{\rm I\!F}_{q}^{*}$ and $m\geq 2$. In paper [5], the authors studied the algebraic structure of $\lambda$-costacyclic ${\rm I\!F}_{q}$-linear additive code over ${\rm I\!F}_{q^{m}}$. In the same paper, they developed a theory for self-orthogonal and self-dual negacyclic ${\rm I\!F}_{q}$-linear codes over ${\rm I\!F}_{q^{l}}$. The authors in paper [26] deduced a theory for additive cyclic and cyclic ${\rm I\!F}_{2}$-linear ACD codes over ${\rm I\!F}_{4}$ of odd length for the trace Euclidean and Hermitian inner product. They also provided a characterization for subfield subcodes. Inspired by these papers, we will study the constacyclic ${\rm I\!F}_{q}$-linear ACP of codes and counting of constacyclic ${\rm I\!F}_{q}$-linear ACP of codes.
@@
The paper is organized as follows. In Section 2, firstly, we sketch basic definitions, notations, and results on ${\rm I\!F}_{q}$-linear additive codes over ${\rm I\!F}_{q^{m}}$, and secondly, we recall some useful results for our context from constacyclic ${\rm I\!F}_{q}$-linear additive codes over ${\rm I\!F}_{q^{m}}$. Section 3 deals with the characterization of ACP of codes with respect to the general inner product on ${\rm I\!F}_{q^{m}}^{n}$. Some constructions of ACP of codes are presented in Section 4. Further, we count a formula for constacyclic ACP of codes in Section 5. Finally, we show an application on EAQEC codes in Section 6 and obtain some examples of EAQEC codes. The paper concludes in Section 7.
@@ 2.1 ${\rm I\!F}_{q}$-linear additive codes
Let ${\rm I\!F}_{q}$ and ${\rm I\!F}_{q^{m}}$ be the finite fields with cardinality $q$ and $q^{m}$, respectively, and characteristic $p$ (see for more details [12]). A nonempty subset $C$ of ${\rm I\!F}_{q^{m}}^{n}$ (where $m>1$) is called an ${\rm I\!F}_{q}$-linear additive code if $C$ is an ${\rm I\!F}_{q}$-linear subspace of ${\rm I\!F}_{q^{m}}^{n}$. Note that an ${\rm I\!F}_{q}$-linear additive code of ${\rm I\!F}_{q^{m}}^{n}$ is an ${\rm I\!F}_{q}$-linear code over ${\rm I\!F}_{q^{m}}$ of length $n$. Now, we define a bilinear mapping
@@
$((a_{1},a_{2},\ldots,a_{n}),(b_{1},b_{2},\ldots,a_{n}))$ $\mapsto\mathcal{B}((a_{1},a_{2},\ldots,a_{n}),(b_{1},b_{2},\ldots,a_{n}))=\sum\limits_{i=1}^{n}Tr(\mu_{i}a_{i}\pi(b_{\sigma(i)}))$ (1)
@@
    $Rad_{L}(B)=\{a\in{\rm I\!F}_{q^{m}}^{n}~{}|~{}\mathcal{B}(a,x)=0~{}\text{for~{}all}~{}x\in{\rm I\!F}_{q^{m}}^{n}\}$ is an ${\rm I\!F}_{q}$-linear subspace of ${\rm I\!F}_{q^{m}}^{n}$;
@@
    $Rad_{R}(B)=\{a\in{\rm I\!F}_{q^{m}}^{n}~{}|~{}\mathcal{B}(y,a)=0~{}\text{for~{}all}~{}y\in{\rm I\!F}_{q^{m}}^{n}\}$ is an ${\rm I\!F}_{q}$-linear subspace of ${\rm I\!F}_{q^{m}}^{n}$.
For an ${\rm I\!F}_{q}$-linear additive code $C$ over ${\rm I\!F}_{q^{m}}$, the left-dual of $C$ is denoted as $C^{\perp_{L}}$, defined by
$C^{\perp_{L}}=\{a\in{\rm I\!F}_{q^{m}}^{n}~{}|~{}\mathcal{B}(a,c)=0~{}\text{for~{}all}~{}c\in C\}$
and the right-dual of $C$ is denoted as $C^{\perp_{R}}$, defined by
$C^{\perp_{R}}=\{a\in{\rm I\!F}_{q^{m}}^{n}~{}|~{}\mathcal{B}(c,a)=0~{}\text{for~{}all}~{}c\in C\}.$
Note that $C^{\perp_{L}}$ and $C^{\perp_{R}}$ are ${\rm I\!F}_{q}$-linear subspace of ${\rm I\!F}_{q^{m}}^{n}$. It is not difficult to show that $C^{\perp_{L}}$ and $C^{\perp_{R}}$ satisfy the conditions
$\dim_{{\rm I\!F}_{q}}(C)+\dim_{{\rm I\!F}_{q}}(C^{\perp_{L}})=nm$ (2)
$\dim_{{\rm I\!F}_{q}}(C)+\dim_{{\rm I\!F}_{q}}(C^{\perp_{R}})=nm$ (3)
$(C^{\perp_{L}})^{\perp_{R}}=(C^{\perp_{R}})^{\perp_{L}}=C.$ (4)
@@ 2.2 Constacyclic ${\rm I\!F}_{q}$-linear additive codes
In this subsection, we assume that $\gcd(n,p)=1$. Suppose that $\lambda$ is an element of ${\rm I\!F}_{q}^{*}$ with multiplicative order $t$. Consequently, $\gcd(nt,p)=1$ as $t$ is divisor of $q-1$. Now we spell out some notations for our contexts as follows.
@@
  • Denote $C_{i}^{(b)}=\{i,ib,ib^{2},\ldots\}\pmod{N}$ as the $b$-cyclotomic coset containing $i$ modulo $N$ where $i\in\{0,1,\ldots,n-1\}$ and $b$ is either $q$ or $q^{m}$.
  • Denote the cardinality of $C_{i}^{(b)}$ as $|C_{i}^{(b)}|$.
We need the following lemma for the enrichment of our context. So, we deliver the lemma as follows.
@@
    If $m$ is a factor of $|C^{(q)}_{i}|$, then $|C_{i}^{(q)}|=m|C_{i}^{(q^{m})}|$ and $C_{i}^{(q)}=C_{i}^{(q^{m})}\cup C_{iq}^{(q^{m})}\cup\ldots\cup C_{iq^{m-1}}^{(q^{m})}$, where $C_{iq^{j}}^{(q^{m})},0\leq j\leq m-1$ are pairwise disjoint with same cardinality.
Let $X$ be an indeterminate over ${\rm I\!F}_{q^{m}}$. Now consider the quotient ring
$\mathcal{R}^{(q)}_{n,\lambda}$ $=\dfrac{{\rm I\!F}_{q}[X]}{(X^{n}-\lambda)}=\bigoplus_{i=0}^{s}\dfrac{{\rm I\!F}_{q}[X]}{(p_{i}(X))}=\bigoplus_{i=0}^{s}\mathcal{K}_{i}$ (5)
where $X^{n}-\lambda=\prod\limits_{i=0}^{s}p_{i}(X)$ with each $p_{i}(X)$ is an irreducible polynomial in ${\rm I\!F}_{q}[X]$, corresponding $q$-cyclotomic coset $C^{(q)}_{1+tj}$ and $\mathcal{K}_{j}=\dfrac{{\rm I\!F}_{q}[X]}{(p_{j}(X))}$ for $0\leq j\leq s$. Further, consider the quotient ring
$\mathcal{R}^{(q^{m})}_{n,\lambda}$ $=\dfrac{{\rm I\!F}_{q^{m}}[X]}{(X^{n}-\lambda)}=\bigoplus_{i=0}^{\rho}\dfrac{{\rm I\!F}_{q^{m}}[X]}{(M_{i}(X))}=\bigoplus_{i=0}^{\rho}\mathcal{I}_{i}$ (6)
@@
Here each $q_{i}(X)$ is an irreducible polynomial in ${\rm I\!F}_{q^{m}}[X]$, corresponding $q^{m}$-cyclotomic coset $C^{(q)}_{1+tj}$ for $0\leq j\leq s$. Denote $\mathcal{F}_{i}=\oplus_{h=1}^{m}\mathcal{I}_{r+(i-1)m+h}$, for $1\leq i\leq s-r$. For more details to see [5, Section $2$].
@@
[5, Lemma 3]
  1. $(a)$
    Each nonempty $\lambda$-constacyclic ${\rm I\!F}_{q}$-linear additive code $C$ in ${\rm I\!F}_{q^{m}}^{n}$ can be uniquely expressed as
    $C=C_{1}\bigoplus C_{2}\bigoplus\cdots\bigoplus C_{s},$ (7)
    where each $C_{i}$ is a unique $\mathcal{K}_{i}$-subspace of $\mathcal{I}_{i}$, for $0\leq i\leq r$ and $C_{r+k}$ is a unique $\mathcal{K}_{i}$-subspace of $\mathcal{F}_{r+k}$ for $1\leq k\leq s-r$.
  2. $(b)$
    Conversely, let $D_{i}$ be an $\mathcal{K}_{i}$-subspace of $\mathcal{I}_{i}$, for $0\leq i\leq r$ and $D_{r+k}$ be an $\mathcal{K}_{i}$-subspace of $\mathcal{F}_{r+k}$ for $1\leq k\leq s-r$. Then the direct sum
    $D=D_{1}\bigoplus D_{2}\bigoplus\cdots\bigoplus D_{s}$ (8)

### v3

@@ Abstract
An additive code is an ${\rm I\!F}_{q}$-linear subspace of ${\rm I\!F}_{q^{m}}^{n}$ over ${\rm I\!F}_{q^{m}}$, which is not a linear subspace over ${\rm I\!F}_{q^{m}}$. Linear complementary pairs (LCP) of codes have important roles in cryptography, such as increasing the speed and capacity of digital communication and strengthening security by improving the encryption necessities to resist cryptanalytic attacks. This paper studies an algebraic structure of additive complementary pairs (ACP) of codes over ${\rm I\!F}_{q^{m}}$. Further, we characterize an ACP of codes in analogous generator matrices and parity check matrices. Additionally, we identify a necessary condition for an ACP of codes. Besides, we present some constructions of an ACP of codes over ${\rm I\!F}_{q^{m}}$ from LCP codes over ${\rm I\!F}_{q^{m}}$ and also from an LCP of codes over ${\rm I\!F}_{q}$. Finally, we study the constacyclic ACP of codes over ${\rm I\!F}_{q^{m}}$ and the counting of the constacyclic ACP of codes.
@@ 1 Introduction
Linear complementary pairs (LCP) of codes, which were introduced by Bhasin et al. in [17], are extensively explored for wide application in cryptography (see [2, 4]). Further, we refer to some papers [1, 5, 9] on the LCP of codes over a finite field, and we point out that complementary pairs of codes without linearity properties have an essential role in constructing quantum codes (see [18]). Thus, it is a useful context for studying non-linear complementary pairs of codes in coding theory. In particular, we study additive complementary pairs (ACP) of codes that extend the analogous of additive complementary dual (ACD) codes. With this point of view, lots of research has been done for additive complementary dual (ACD) codes (see [19, 20, 21]). LCP of codes is a generalization of linear complementary dual (LCD) codes. The notion of LCD codes was introduced by Massy in [15] and later studied in [6, 7, 8, 9, 16].
@@
Besides, the study of additive codes has gathered significant attention due to their theoretical significance and practical applications. One prominent application of additive codes lies in quantum error correction, where they are employed to protect quantum information from errors induced by noise or other environmental factors. Additionally, additive codes find utility in secret-sharing schemes, where they facilitate the secure distribution of confidential information among multiple parties, ensuring that only authorized subsets of participants can reconstruct the original secret [14]. Additive codes represent a crucial class of codes within coding theory, introduced by Delsarte et al. [11]. Generally, additive codes are subgroups of the underlying abelian group. Further, Huffman [13] provided an algebraic structure for additive cyclic codes over ${\rm I\!F}_{4}$. Through ongoing research and exploration, the potential of additive codes continues to expand, offering valuable insights and solutions in the realms of information theory, quantum computing, and cryptography (see [12]). Recently, Shi et al. [20] developed a theory of ACD codes over ${\rm I\!F}_{4}$ for trace Euclidean and Hermitian inner products. The authors introduced a nice construction of ACP codes over ${\rm I\!F}_{4}$ from binary codes in the same paper. They also established that ACD codes are potentially related to LCD codes. In the same spirit, Choi et al. [10] studied ACP codes over a finite field ${\rm I\!F}_{q^{m}}$ with $m\geq 2$ in a general setup of trace inner product. They also provided new construction of ACP codes with respect to trace-Euclidean, Hermitian, and Galois inner products. This general setup showed that ACD codes are related to LCD codes. In this same paper, they further computed good numerical examples of ACP codes. Motivated by these papers, we study the theory of additive complementary pair (ACP) of codes. We further derive some construction of ACP of codes. This paper establishes a relationship between the LCP of codes and the ACP of codes.
@@
On the other hand, an $\dfrac{{\rm I\!F}_{q}[X]}{\left(X^{n}-\lambda\right)}$-submodule of $\dfrac{{\rm I\!F}_{q^{m}}[X]}{\left(X^{n}-\lambda\right)}$ is a $\lambda$-constacyclic ${\rm I\!F}_{q}$-linear additive code over ${\rm I\!F}_{q^{m}}$, where $\lambda\in{\rm I\!F}_{q}^{*}$ and $m\geq 2$. In paper [3], the authors studied the algebraic structure of $\lambda$-constacyclic ${\rm I\!F}_{q}$-linear additive code over ${\rm I\!F}_{q^{m}}$. In the same paper, they developed a theory for self-orthogonal and self-dual negacyclic ${\rm I\!F}_{q}$-linear codes over ${\rm I\!F}_{q^{l}}$. The authors in paper [21] deduced a theory for additive cyclic and cyclic ${\rm I\!F}_{2}$-linear ACD codes over ${\rm I\!F}_{4}$ of odd length for the trace Euclidean and Hermitian inner product. They also provided a characterization for subfield subcodes. Inspired by these papers, we will study the constacyclic ${\rm I\!F}_{q}$-linear ACP of codes and counting of constacyclic ${\rm I\!F}_{q}$-linear ACP of codes.
@@
The paper is organized as follows. In Section 2, firstly, we sketch basic definitions, notations, and results on ${\rm I\!F}_{q}$-linear additive codes over ${\rm I\!F}_{q^{m}}$, and secondly, we recall some useful results for our context from constacyclic ${\rm I\!F}_{q}$-linear additive codes over ${\rm I\!F}_{q^{m}}$. Section 3 deals with the characterization of ACP of codes with respect to the general inner product on ${\rm I\!F}_{q^{m}}^{n}$. Some constructions of ACP of codes are presented in Section 4. Further, we count a formula for constacyclic ACP of codes in Section 5. Finally, the paper concludes in Section 6.
@@ 2.1 ${\rm I\!F}_{q}$-linear additive codes
Let ${\rm I\!F}_{q}$ and ${\rm I\!F}_{q^{m}}$ be the finite fields with cardinality $q$ and $q^{m}$, respectively, and characteristic $p$. A nonempty subset $\mathcal{C}$ of ${\rm I\!F}_{q^{m}}^{n}$ (where $m>1$) is called an ${\rm I\!F}_{q}$-linear additive code if $\mathcal{C}$ is an ${\rm I\!F}_{q}$-linear subspace of ${\rm I\!F}_{q^{m}}^{n}$ (see for more details [10]). Note that an ${\rm I\!F}_{q}$-linear additive code of ${\rm I\!F}_{q^{m}}^{n}$ is an ${\rm I\!F}_{q}$-linear code over ${\rm I\!F}_{q^{m}}$ of length $n$. Now, we define a bilinear mapping
@@
$((a_{1},a_{2},\ldots,a_{n}),(b_{1},b_{2},\ldots,b_{n}))$ $\mapsto\mathcal{B}((a_{1},a_{2},\ldots,a_{n}),(b_{1},b_{2},\ldots,b_{n}))=\sum\limits_{i=1}^{n}Tr(\mu_{i}a_{i}\pi(b_{\sigma(i)}))$ (1)
@@
    $Rad_{L}(B)=\{\mathbf{a}\in{\rm I\!F}_{q^{m}}^{n}~{}|~{}\mathcal{B}(\mathbf{a},\mathbf{x})=0~{}\text{for~{}all}~{}\mathbf{x}\in{\rm I\!F}_{q^{m}}^{n}\}$ is an ${\rm I\!F}_{q}$-linear subspace of ${\rm I\!F}_{q^{m}}^{n}$;
@@
    $Rad_{R}(B)=\{\mathbf{a}\in{\rm I\!F}_{q^{m}}^{n}~{}|~{}\mathcal{B}(\mathbf{y},\mathbf{a})=0~{}\text{for~{}all}~{}\mathbf{y}\in{\rm I\!F}_{q^{m}}^{n}\}$ is an ${\rm I\!F}_{q}$-linear subspace of ${\rm I\!F}_{q^{m}}^{n}$.
For an ${\rm I\!F}_{q}$-linear additive code $\mathcal{C}$ over ${\rm I\!F}_{q^{m}}$, the left-dual of $\mathcal{C}$ is denoted as $\mathcal{C}^{\perp_{L}}$, defined by
$\mathcal{C}^{\perp_{L}}=\{\mathbf{a}\in{\rm I\!F}_{q^{m}}^{n}~{}|~{}\mathcal{B}(\mathbf{a},\mathbf{c})=0~{}\text{for~{}all}~{}\mathbf{c}\in\mathcal{C}\}$
and the right-dual of $\mathcal{C}$ is denoted as $\mathcal{C}^{\perp_{R}}$, defined by
$\mathcal{C}^{\perp_{R}}=\{\mathbf{a}\in{\rm I\!F}_{q^{m}}^{n}~{}|~{}\mathcal{B}(\mathbf{c},\mathbf{a})=0~{}\text{for~{}all}~{}\mathbf{c}\in\mathcal{C}\}.$
Note that $\mathcal{C}^{\perp_{L}}$ and $\mathcal{C}^{\perp_{R}}$ are ${\rm I\!F}_{q}$-linear subspace of ${\rm I\!F}_{q^{m}}^{n}$. It is not difficult to show that $\mathcal{C}^{\perp_{L}}$ and $\mathcal{C}^{\perp_{R}}$ satisfy the conditions (for more details, we refer [10])
$\dim_{{\rm I\!F}_{q}}(\mathcal{C})+\dim_{{\rm I\!F}_{q}}(\mathcal{C}^{\perp_{L}})=nm$ (2)
$\dim_{{\rm I\!F}_{q}}(\mathcal{C})+\dim_{{\rm I\!F}_{q}}(\mathcal{C}^{\perp_{R}})=nm$ (3)
$(\mathcal{C}^{\perp_{L}})^{\perp_{R}}=(\mathcal{C}^{\perp_{R}})^{\perp_{L}}=\mathcal{C}.$ (4)
@@ 2.2 Constacyclic ${\rm I\!F}_{q}$-linear additive codes
In this subsection, we assume that $\gcd(n,p)=1$. Suppose that $\lambda\in{\rm I\!F}_{q}^{*}$ with multiplicative order $t$. Consequently, $\gcd(nt,p)=1$ as $t$ is divisor of $q-1$. Now, we spell out some notations for our contexts as follows.
@@
  • Denote $\mathcal{C}_{i}^{(b)}=\{i,ib,ib^{2},\ldots\}\pmod{N}$ as the $b$-cyclotomic coset containing $i$ modulo $N$ where $i\in\{0,1,\ldots,n-1\}$ and $b$ is either $q$ or $q^{m}$.
  • Denote the cardinality of $\mathcal{C}_{i}^{(b)}$ as $|\mathcal{C}_{i}^{(b)}|$.
We following lemma is presented to enrich our context.
@@
    If $m$ is a factor of $|\mathcal{C}^{(q)}_{i}|$, then $|\mathcal{C}_{i}^{(q)}|=m|\mathcal{C}_{i}^{(q^{m})}|$ and $\mathcal{C}_{i}^{(q)}=\mathcal{C}_{i}^{(q^{m})}\cup\mathcal{C}_{iq}^{(q^{m})}\cup\ldots\cup\mathcal{C}_{iq^{m-1}}^{(q^{m})}$, where $\mathcal{C}_{iq^{j}}^{(q^{m})},0\leq j\leq m-1$ are pairwise disjoint with same cardinality.
Let $X$ be an indeterminate over ${\rm I\!F}_{q^{m}}$. Consider the quotient ring
$\mathcal{R}^{(q)}_{n,\lambda}$ $=\dfrac{{\rm I\!F}_{q}[X]}{(X^{n}-\lambda)}=\bigoplus_{i=0}^{s}\dfrac{{\rm I\!F}_{q}[X]}{(p_{i}(X))}=\bigoplus_{i=0}^{s}\mathcal{K}_{i}$
where $X^{n}-\lambda=\prod\limits_{i=0}^{s}p_{i}(X)$ with each $p_{i}(X)$ is an irreducible polynomial in ${\rm I\!F}_{q}[X]$, corresponding $q$-cyclotomic coset $\mathcal{C}^{(q)}_{1+tj}$ and $\mathcal{K}_{j}=\dfrac{{\rm I\!F}_{q}[X]}{(p_{j}(X))}$ for $0\leq j\leq s$. Further, consider the quotient ring
$\mathcal{R}^{(q^{m})}_{n,\lambda}$ $=\dfrac{{\rm I\!F}_{q^{m}}[X]}{(X^{n}-\lambda)}=\bigoplus_{i=0}^{\rho}\dfrac{{\rm I\!F}_{q^{m}}[X]}{(M_{i}(X))}=\bigoplus_{i=0}^{\rho}\mathcal{I}_{i}$
@@
Here each $q_{i}(X)$ is an irreducible polynomial in ${\rm I\!F}_{q^{m}}[X]$, corresponding $q^{m}$-cyclotomic coset $\mathcal{C}^{(q)}_{1+tj}$ for $0\leq j\leq s$. Denote $\mathcal{F}_{i}=\oplus_{h=1}^{m}\mathcal{I}_{r+(i-1)m+h}$, for $1\leq i\leq s-r$. For more details, see [3, Section $2$].
@@
  1. 1)1)1 )
    Each nonempty $\lambda$-constacyclic ${\rm I\!F}_{q}$-linear additive code $\mathcal{C}$ in ${\rm I\!F}_{q^{m}}^{n}$ can be uniquely expressed as
    $\mathcal{C}=\mathcal{C}_{1}\bigoplus\mathcal{C}_{2}\bigoplus\cdots\bigoplus\mathcal{C}_{s},$ (5)
    where each $\mathcal{C}_{i}$ is a unique $\mathcal{K}_{i}$-subspace of $\mathcal{I}_{i}$, for $0\leq i\leq r$ and $\mathcal{C}_{r+k}$ is a unique $\mathcal{K}_{i}$-subspace of $\mathcal{F}_{r+k}$ for $1\leq k\leq s-r$.
  2. 2)2)2 )
    Conversely, let $\mathcal{D}_{i}$ be an $\mathcal{K}_{i}$-subspace of $\mathcal{I}_{i}$, for $0\leq i\leq r$ and $\mathcal{D}_{r+k}$ be an $\mathcal{K}_{i}$-subspace of $\mathcal{F}_{r+k}$ for $1\leq k\leq s-r$. Then, the direct sum
    $\mathcal{D}=\mathcal{D}_{1}\bigoplus\mathcal{D}_{2}\bigoplus\cdots\bigoplus\mathcal{D}_{s}$ (6)
@@
Proof.
The proof of Proposition $1$ follows from [3, Lemma 3.1], and the proof of Proposition $2$ is a straightforward exercise. ∎